Ask your own question, for FREE!
Mathematics 16 Online
OpenStudy (anonymous):

Suppose that z~ N(0,1). a) Find the prob that z < -1.96 b) Find the prob that z => 1.96 c) Find theprob that 0 < z < 1.645

OpenStudy (mathmate):

N(0,1) means that z follows the normal distribution, with parameters standard deviation 1 and mean 0. The table of normal distribution is available at many sources, such as your calculator, or the following link: http://www.math.unb.ca/~knight/utility/NormTble.htm Since the area under a probability distribution equals 1.0 and the distribution is symmetrical about the mean, only half of the table (betwen z=0 to infinity) needs to be tabulated. To look for the probability of a negative z-value, you can look up the positive value and subtract from 1.0. Example P(-0.5)=1-P(0.5)=1-0.6915=0.3085

Can't find your answer? Make a FREE account and ask your own questions, OR help others and earn volunteer hours!

Join our real-time social learning platform and learn together with your friends!
Latest Questions
Arriyanalol: i need help
1 day ago 7 Replies 1 Medal
Arriyanalol: CONGRATS to me
1 day ago 8 Replies 2 Medals
Legomaster346: What colleges are good that are near Arab Alabama?
1 week ago 5 Replies 0 Medals
Arriyanalol: i need help
1 week ago 6 Replies 2 Medals
Can't find your answer? Make a FREE account and ask your own questions, OR help others and earn volunteer hours!

Join our real-time social learning platform and learn together with your friends!